OpenStudy (anonymous):

The frequency of a particular radio wave is approximately 3.00 x 10^5 Hz. The speed is 3.00 x 10^8 m/s. Calculate the wavelength, and remember to include the unit, m.

OpenStudy (anonymous):

speed = frequency * distance travelled 3*10^8 = 3* 10^5 * d d = 3*10^8/(3*10^5) d = 10^3 m or 1km